# ZTF SN Ia DR2: Overview

### v1

@@ Acknowledgements.
- Ruppin et al. (2024, ZTFDR2) Ruppin, F., Rigault, M., Ginolin, M., et al. 2024, arXiv:2406.01108. doi:10.48550/arXiv.2406.01108
- Scolnic et al. (2018) Scolnic, D. M., Jones, D. O., Rest, A., et al. 2018, ApJ, 859, 101. doi:10.3847/1538-4357/aab9bb
- Scolnic et al. (2022) Scolnic, D., Brout, D., Carr, A., et al. 2022, ApJ, 938, 113. doi:10.3847/1538-4357/ac8b7a
- Silverman et al. (2015) Silverman, J. M., Vinkó, J., Marion, G. H., et al. 2015, MNRAS, 451, 1973. doi:10.1093/mnras/stv1011
- Smartt et al. (2015) Smartt, S. J., Valenti, S., Fraser, M., et al. 2015, A&A, 579, A40. doi:10.1051/0004-6361/201425237
- Smith et al. (2012) Smith, M., Nichol, R. C., Dilday, B., et al. 2012, ApJ, 755, 61. doi:10.1088/0004-637X/755/1/61
- Smith et al. (2020) Smith, M., Sullivan, M., Wiseman, P., et al. 2020, MNRAS, 494, 4426. doi:10.1093/mnras/staa946
@@
- Spergel et al. (2003) Spergel, D. N., Verde, L., Peiris, H. V., et al. 2003, ApJS, 148, 175. doi:10.1086/377226
- Steele et al. (2004) Steele, I. A., Smith, R. J., Rees, P. C., et al. 2004, Proc. SPIE, 5489, 679. doi:10.1117/12.551456
- Sullivan et al. (2006) Sullivan, M., Le Borgne, D., Pritchet, C. J., et al. 2006, ApJ, 648, 868. doi:10.1086/506137
- Sullivan et al. (2010) Sullivan, M., Conley, A., Howell, D. A., et al. 2010, MNRAS, 406, 782. doi:10.1111/j.1365-2966.2010.16731.x
- Taylor et al. (2021) Taylor, G., Lidman, C., Tucker, B. E., et al. 2021, MNRAS, 504, 4111. doi:10.1093/mnras/stab962
@@
- Tucker et al. (2020) Tucker, M. A., Shappee, B. J., Vallely, P. J., et al. 2020, MNRAS, 493, 1044. doi:10.1093/mnras/stz3390
- van der Walt et al., (2019) van der Walt et al., (2019) JOSS, 4(37), 1247, doi.org/10.21105/joss.01247
- Vincenzi et al. (2024) Vincenzi, M., Brout, D., Armstrong, P., et al. 2024, arXiv:2401.02945. doi:10.48550/arXiv.2401.02945
- Whelan & Iben (1973) Whelan, J. & Iben, I. 1973, ApJ, 186, 1007. doi:10.1086/152565
- Wiseman et al. (2022) Wiseman, P., Vincenzi, M., Sullivan, M., et al. 2022, MNRAS, 515, 4587. doi:10.1093/mnras/stac1984
- Yao et al. (2019) Yao, Y., Miller, A. A., Kulkarni, S. R., et al. 2019, ApJ, 886, 152. doi:10.3847/1538-4357/ab4cf5
- Yaron & Gal-Yam (2012) Yaron, O. & Gal-Yam, A. 2012, PASP, 124, 668. doi:10.1086/666656

### v2

@@ Acknowledgements.
- Ruppin et al. (2024) Ruppin, F., Rigault, M., Ginolin, M., et al. 2024, arXiv:2406.01108. doi:10.48550/arXiv.2406.01108, (ZTFSI)
- Senzel et al. (2024) Senzel, R., Maguire, K., Burgaz, U., et al. 2024, arXiv:2411.11986. doi:10.48550/arXiv.2411.11986 (ZTFSI)
- Scolnic et al. (2018) Scolnic, D. M., Jones, D. O., Rest, A., et al. 2018, ApJ, 859, 101.
- Scolnic et al. (2022) Scolnic, D., Brout, D., Carr, A., et al. 2022, ApJ, 938, 113.
- Silverman et al. (2015) Silverman, J. M., Vinkó, J., Marion, G. H., et al. 2015, MNRAS, 451, 1973.
- Smartt et al. (2015) Smartt, S. J., Valenti, S., Fraser, M., et al. 2015, A&A, 579, A40.
- Smith et al. (2012) Smith, M., Nichol, R. C., Dilday, B., et al. 2012, ApJ, 755, 61.
- Smith et al. (2020) Smith, M., Sullivan, M., Wiseman, P., et al. 2020, MNRAS, 494, 4426.
@@
- Spergel et al. (2003) Spergel, D. N., Verde, L., Peiris, H. V., et al. 2003, ApJS, 148, 175.
- Steele et al. (2004) Steele, I. A., Smith, R. J., Rees, P. C., et al. 2004, Proc. SPIE, 5489, 679.
- Sullivan et al. (2006) Sullivan, M., Le Borgne, D., Pritchet, C. J., et al. 2006, ApJ, 648, 868.
- Sullivan et al. (2010) Sullivan, M., Conley, A., Howell, D. A., et al. 2010, MNRAS, 406, 782.
- Taylor et al. (2021) Taylor, G., Lidman, C., Tucker, B. E., et al. 2021, MNRAS, 504, 4111.
- Terwel et al. (2024) Terwel, J. H., Maguire, K., Dimitriadis, G., et al. 2024, arXiv:2402.16962. doi:10.48550/arXiv.2402.16962, (ZTF SI)
@@
- Tucker et al. (2020) Tucker, M. A., Shappee, B. J., Vallely, P. J., et al. 2020, MNRAS, 493, 1044.
- van der Walt et al., (2019) van der Walt et al., (2019) JOSS, 4(37), 1247,
- Vincenzi et al. (2024) Vincenzi, M., Brout, D., Armstrong, P., et al. 2024, arXiv:2401.02945.
- Whelan & Iben (1973) Whelan, J. & Iben, I. 1973, ApJ, 186, 1007.
- Wiseman et al. (2022) Wiseman, P., Vincenzi, M., Sullivan, M., et al. 2022, MNRAS, 515, 4587.
- Yao et al. (2019) Yao, Y., Miller, A. A., Kulkarni, S. R., et al. 2019, ApJ, 886, 152.
- Yaron & Gal-Yam (2012) Yaron, O. & Gal-Yam, A. 2012, PASP, 124, 668.